News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 23 June 2000 7.00am EDT

Please note: this report will now be distributed each Friday.

During the last week, Chandra completed the observing schedule as planned.

The HETG was retracted manually as planned upon the completion of a series of grating observations at 10.30pm EDT on June 19. The retraction was commanded using hardware commands with a 155 second timed retraction followed by a 1 second short retraction. Telemetry was consistent with the grating having stopped just short of the bumper after the first move and slightly compressing the bumper without tripping the limit switch on the second short move. A real-time insertion and retraction are planned for 8:15pm EDT on June 25 and 10.25pm EDT on June 26 respectively. Further HETG moves are planned no earlier than the July 16 load and LETG moves no earlier than August. A technical meeting to discuss operational and software patch options will be held at the OCC on June 27.

A routine real-time procedure was completed on 5/20 to dump the CPE warm start counter and showed no change in value (1).

The planned observations for the coming week are as follows.

------------------------------------------
PKS 1830-211           ACIS-I/HETG     Jun 26 (manual insert/retract)
SS 433 (East Jet)      ACIS-S
Q2345+007              ACIS-S          Jun 27
NGC 1291               ACIS-S
Pictor A               ACIS-S          Jun 28
Radiation Belts
A907                   ACIS-I
1156+295               ACIS-I          Jun 29
NGC 4565               ACIS-S
NGC 6121               ACIS-S          Jun 30
Kepler SNR             ACIS-S
Radiation Belts
GRO J1655-40           ACIS-S          Jul  1
HCG 90                 ACIS-I          Jul  2
M31 Transient          ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports